OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2 (Points : 1) y – 9 = 2(x – 5) y – 5 = 2(x + 9) y + 9 = 2(x – 5) y – 9 = 2(x + 5) HELP PLEASE

OpenStudy (redheadangel):

so for the first one y-9=5(x-9) and the 2nd y-3=6(x-(-8) y-3=6(x+8)

OpenStudy (redheadangel):

point slope form is y-y1=m(x-x1) the point becomes (x1,y1) and m=slope

OpenStudy (catlover5925):

ok so point slope is (Y - y1) = M(x - x1) M is the slope and plug everything in (Y - 9) = 2(x - 5)

OpenStudy (anonymous):

Thanks for the help @Catlover5925

OpenStudy (catlover5925):

no problem :)